a partially destroyed laboratory data, only the equations giving the two lines of regression of x on y and y on x are available and are respectively, 7x-16y + 9 = 0, 5y - 4x-3 = 0. Calculate


the coefficient of correlation, x and y.


1
Expert's answer
2021-01-26T01:06:47-0500

The given two equation are "7x-16y+9=0" ......(1) and "5y-4x-3=0" ........(2)

Suppose "7x-16y+9=0" is regression equation of "y" on "x"

"\implies 7x-16y+9=0"

"\implies 16y=7x+9"

"\implies y=\frac{7}{16}x+\frac{9}{16}"

"\implies y=0.4375 x+0.5625"

"\therefore byx=0.4375"

Again suppose "5y-4x-3=0" is regression equation of "x" on "y"

"\implies 5y-4x-3=0"

"\implies 4x=5y-3"

"\implies x=\frac{5}{4}y-\frac{3}{4}"

"\implies x=1.25y-0.75"

"\therefore byx=1.25"

So the coefficient of correlation "x" and "y" is ,

"r=+\sqrt{byx.bxy}" [As "byx>0" and "bxy>0]"

"\therefore r=\sqrt{\frac{7}{16}.\frac{5}{4}}=\sqrt{0.546875}=0.7395"